OpenStudy (anonymous):

a minivan sells for $32000 plus gst and pst. the dealership predicts that in 3 yrs this price will increase 15%. how much should you incest today at 7.25% per annum, compounded semi-annually, to buy the new model in 3 years?

OpenStudy (anonymous):

In 3 years, the vehicle will cost 1.15 (32000) = 36800 Interest compounded ½-yearly @ 7.25% means that, over 3 years, there will be 6 periods at 3.625%, so P (1.03625)^6 = 36800 1.23819 P = 36800 P = 36800 / 1.23819 = 29720.8 or $29721 to the nearest dollar.

OpenStudy (anonymous):

i need to find how much i should invest to pay off the price of the new model... if the price in 3 years is 38000, don't i need to use the equation \[PV= A / (1 + i )^{n}\]

OpenStudy (anonymous):

The compound interest equation: A = P ( 1 + r/n ) ^ (nt) 36,800 = P ( 1 + .0725/2 ) ^6 => P = 36,800 / ( 1 + .0725/2 ) ^6 = $29.721